OpenStudy (anonymous):

how do I prove 2tanx/sinx=cscxtanx+secx

OpenStudy (anonymous):

tan(x)/sin(x)= sec(x)

OpenStudy (anonymous):

so this is 2 sec(x)

OpenStudy (anonymous):

so we have to prove csc x tan(x) = sec(x) 1/sin(x) * sin(x)/cos(x)= 1/cos(x)= sec(x)

OpenStudy (anonymous):

How did you get tan(x)/sin(x)= sec(x)

OpenStudy (anonymous):

tan(x)= sin(x)/cos x sin x / cos x * 1/ sin(x) sin(x) cancels , leaving 1/cos x = sec x

OpenStudy (anonymous):

oh ok

OpenStudy (anonymous):

oh I see the answer. thank you.

OpenStudy (anonymous):

wait, wouldn't sinx / cos x be multipy by sinx not 1/sinx

OpenStudy (anonymous):

why

OpenStudy (anonymous):

because the sinx is on the denominator and when you do the reciprocial 1/sinx it's sinx

OpenStudy (anonymous):

because the sinx is on the denominator and when you do the it is reciprocial 1/sinx fixed

OpenStudy (anonymous):

fixed? what?

OpenStudy (anonymous):

divide is same as multiplication by it reciprocal so we dividing by sin(x) that's same as multiplying by 1/sin x

OpenStudy (anonymous):

oh ok. I understand

OpenStudy (anonymous):

thanks for your help.

OpenStudy (anonymous):

you are welcome